(c) Use the gradient-y-intercept method to draw the graph of 5x + 2y + 3 = 0 and read the following from your graph: y when x = -1½ and x when
y = ½.

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

Use the gradient-y-intercept method to draw the graph of 5x + 2y + 3 = 0 and read the following from your graph: y when x = -1½ and x when y = ½.
------------------
y = -(5x+3)/2
---
y = (-5/2)x - (3/2)
----------------------
f(-3/2) = (-5/2)(-3/2) - (3/2) = +15/4 - (6/4) = 9/4
---------------
f(1/2) = (-5/2)(1/2) - (3/2) = -5/4 - 6/4 = -11/4
-------------
Cheers,
Stan H.
-----------

Answer by ewatrrr(24785)

5x + 2y + 3 = 0
y = (-5/2)x - 3/2
P(1,-4) and P(0,-3/2) on this line: Plot and connect
(-3/2 ,2) and (-4/5, 1/5) are the points in question